OpenStudy (anonymous):

help please solve for U -4u+7=-10u+37

OpenStudy (solomonzelman):

1) add \(\normalsize\color{black}{ \rm 4u }\) to both sides. 2) subtract \(\normalsize\color{black}{ \rm 37 }\) from both sides. 3) divide by \(\normalsize\color{black}{ \rm -6 }\) on both sides.

OpenStudy (jackofalltradez):

Welp, you just said exactly what I was gonna type, so, medal for you Solomon :)

OpenStudy (anonymous):

so it would be -4u +7 = -10u +37 +4u +4u 7 = -6 +37 +37 +37 44 = -6 44/-6 -6/-6 -7.3333333

OpenStudy (anonymous):

is this write or wrong

OpenStudy (anonymous):

@JackofallTradez

OpenStudy (anonymous):

@SolomonZelman